How long would it take for a ball dropped from the top of a
196-foot building to hit the ground? Round your answer to two decimal places.

Answer by addingup(3677)

To answer your question, we have to use the formula to approximate the distance in feet s that the ball (or any other object) falls freely from rest in t seconds:
s = 16t^2
Now, you may wonder why 16, why not any other number. It's because the instantaneous acceleration, the acceleration of gravity, is 32 feet per second. The coefficient of the t^2 term is HALF the acceleration. When you get to calculus you'll learn how this works.
So here we go:
16t^2 = 196
t^2 = 196/16
t^2 = 12.25
t = sqrt(12.25)
t = 3.5 seconds